i^+j ^+k^and i^+ j^ are unit vectors then angle between them is

Grade:12th pass

1 Answers

Chirag Vaswani
6 years ago
a.b=|a||b| cos(x) where x is the angle between them.
a.b = 1*1+1*1+1*0
      = 2
|a| = ( 1^2+1^2+1^2)^1/2
     = 3^1/2
|b| = 2^1/2
 
hence a.b/|a||b| = 2/2^1/2*3^1/2
                          = (2/3)^1/2
 
therefore x= cos inverse((2/3)^1/2)